find the standard equation of the circle C(3,-2) ; r=4
1
Expert's answer
2019-06-13T08:41:09-0400

In case the center of a circle C(a;b) and the radius r, the standard form of a circle equation is (x-a)2+(y-b)2=r2.

We get the equation of the circle in the form (x-3)2+(y+2)2=42 or (x-3)2+(y+2)2=16.